Explain what tanθ= -3/4 means. include and English explanation.
Thanks,
paulaficarella@gmail.com

Answer by stanbon(75887)

Explain what tanθ= -3/4 means. include and English explanation.
------
Using a unit circle model. tan(theta) = y/x = -3/4 implies
that theta is an angle in QII or QIV.
If theta is in QII x = -4 and y = 3.
If theta is in QIV x = 3 and y = -4.
------------------------------------
Cheers,
Stan H.
=======================